# Towards Understanding Why FixMatch Generalizes Better Than Supervised Learning

### v1

@@ E.2.3 Noise Correlation is Small
Suppose Parameter 10 holds and suppose Induction Hypothesis 15 holds for all iteration $<t$ starting from $T$. For every $l\in[2]$, for every $r\in[m]$, for every $(X,y)\in\tilde{\mathcal{Z}}_{u}$ and $i\in[k]$:
@@
$\langle w_{i,r}^{(t+1)},\xi_{p}\rangle=\langle w_{i,r}^{(t)},\xi_{p}\rangle+\tilde{\Theta}(\frac{\eta}{\tilde{N}_{u}})\overline{\operatorname{ReLU}}^{\prime}(\langle w_{i,r}^{(t)},x_{p}\rangle)(1-\mathbf{logit}_{i}(F^{(t)},X))\pm\frac{\eta}{\sqrt{d}}.$
@@
$\langle w_{i,r}^{(t+1)},\xi_{p}\rangle=\langle w_{i,r}^{(t)},\xi_{p}\rangle-\tilde{\Theta}(\frac{\eta}{\tilde{N}_{u}})\overline{\operatorname{ReLU}}^{\prime}(\langle w_{i,r}^{(t)},x_{p}\rangle)\mathbf{logit}_{i}(F^{(t)},X)\pm\frac{\eta}{\sqrt{d}}.$
Suppose that it satisfies that $|\langle w_{i,r}^{(t)},x_{p}\rangle|\leq A$ for every $t<t_{0}$ where $t_{0}$ is any iteration $T\leq t_{0}\leq T+T^{\prime}$. When $T\leq t\leq T+T_{0}$, we have that
$\langle w_{i,r}^{(t)},\xi_{p}\rangle$ $\leq\langle w_{i,r}^{(T)},\xi_{p}\rangle+\tilde{O}\left(\frac{T_{0}\eta A^{q-1}}{\tilde{N}_{u}}\right)+\frac{T_{0}\eta}{\sqrt{d}}$
$\leq\tilde{o}(\sigma_{0})+\tilde{O}\left(\frac{kA^{q-1}}{\tilde{N}_{u}\sigma_{0}^{q-2}}\right)+\frac{T_{0}\eta}{\sqrt{d}},$
where the last step is because $T_{0}=\tilde{\Theta}(\frac{k}{\eta\sigma_{0}^{q-2}})$. When $T+T_{0}\leq t\leq T+T^{\prime}$, for multi-view data $(X,y)\sim\tilde{\mathcal{Z}}_{u,m}$, based on (15) in Claim 28, we can obtain that
$\langle w_{i,r}^{(t)},\xi_{p}\rangle$ $\leq\langle w_{i,r}^{(T+T_{0})},\xi_{p}\rangle+\tilde{O}\left(\frac{k^{5}A^{q-1}}{s^{2}\tilde{N}_{u}}\right)+\frac{(T^{\prime}-T_{0})\eta}{\sqrt{d}}$
$\leq\tilde{O}\left(\frac{kA^{q-1}}{\tilde{N}_{u}\sigma_{0}^{q-2}}+\frac{k^{5}A^{q-1}}{s^{2}\tilde{N}_{u}}\right)+\frac{T^{\prime}\eta}{\sqrt{d}}.$
@@
$\langle w_{i,r}^{(t)},\xi_{p}\rangle$ $\leq\langle w_{i,r}^{(T+T_{0})},\xi_{p}\rangle+\tilde{O}\left(\frac{T^{\prime}\eta A^{q-1}}{\tilde{N}_{u}}\right)+\frac{(T^{\prime}-T_{0})\eta}{\sqrt{d}}$
$\leq\tilde{O}\left(\frac{kA^{q-1}}{\tilde{N}_{u}\sigma_{0}^{q-2}}+\frac{T^{\prime}\eta A^{q-1}}{\tilde{N}_{u}}\right)+\frac{T^{\prime}\eta}{\sqrt{d}}.$
When $p\in\mathcal{P}_{v_{i,l}}(X)$, we have $|\langle w_{i,r}^{(t)},x_{p}\rangle|\leq\tilde{O}(1)$ from Induction Hypothesis 15. Then plugging in $A=\tilde{O}(1),\tilde{N}_{u}\geq\tilde{\Omega}\left(\frac{k}{\sigma_{0}^{q-1}}\right)$, $\tilde{N}_{u}\geq\tilde{\Omega}\left(\frac{k^{5}}{\sigma_{0}}\right)$ and $\tilde{N}_{u}\geq\eta T^{\prime}\mathrm{poly}(k)$, we can obtain that $|\langle w_{i,r}^{(t)},\xi_{p}\rangle|\leq\tilde{o}(\sigma_{0})$.
When $p\in\mathcal{P}(X)\setminus(\mathcal{P}_{v_{i,1}}(X)\cup\mathcal{P}_{v_{i,2}}(X))$, we have $|\langle w_{i,r}^{(t)},x_{p}\rangle|\leq\tilde{O}(\sigma_{0})$ from the Induction Hypothesis 15. Then plugging in $A=\tilde{O}(\sigma_{0}),\tilde{N}_{u}\geq k^{5}$ and $\tilde{N}_{u}\geq\eta T^{\prime}\mathrm{poly}(k)$, we can obtain that $|\langle w_{i,r}^{(t)},\xi_{p}\rangle|\leq\tilde{O}(\sigma_{0})$.
When $p\in[P]\setminus\mathcal{P}(X)$, we have $|\langle w_{i,r}^{(t)},x_{p}\rangle|\leq\tilde{O}(\sigma_{0}\gamma k)$ from Induction Hypothesis 15. Then plugging in $A=\tilde{O}(\sigma_{0}\gamma k),\tilde{N}_{u}\geq k^{5}$ and $\tilde{N}_{u}\geq\eta T^{\prime}\mathrm{poly}(k)$, we can obtain that $|\langle w_{i,r}^{(t)},\xi_{p}\rangle|\leq\tilde{O}(\sigma_{0}\gamma k)$.
@@ E.2.4 Diagonal Correlations are Nearly Non-Negative
Suppose Parameter 10 holds and suppose Induction Hypothesis 15 holds for all iteration $<t$ starting from $T$. Then,
@@
From Lemma D.27 in Allen-Zhu & Li (2023), we know $\langle w_{i,r}^{(t)},v_{i,l}\rangle\geq-\tilde{O}(\sigma_{0})$ for every iteration $t\leq T$. Now we consider any iteration $t>T$ so that $\langle w_{i,r}^{(t)},v_{i,l}\rangle\leq-\widetilde{\Omega}(\sigma_{0})$. We start from this iteration to see how negative the next iterations can be. Without loss of generality, we consider the case when $\langle w_{i,r}^{(t^{\prime})},v_{i,l}\rangle\leq-\widetilde{\Omega}(\sigma_{0})$ holds for every $t^{\prime}\geq t$. By Claim 19 and Claim 20,
@@
We first consider every $t\leq T+T_{0}$. Using Claim 23 we have $\mathbf{logit}_{i}(F^{(t)},X)=O\left(\frac{1}{k}\right)$, which implies
@@
As for $t>T+T_{0}$, combining with Claim 26 and the fact that $\mathbf{logit}_{i}(F^{(t)},X)\leq 1-\mathbf{logit}_{y}(F^{(t)},X)$ for $i\neq y$, we have
@@ E.2.5 Proof of Induction Hypothesis 15
Under Parameter Assumption 10, for any $m=\mathrm{polylog}(k)$ and sufficiently small $\eta\leq\frac{1}{\mathrm{poly}(k)}$, our Induction Hypothesis 15 holds for all iterations $t=T,T+1,\dots,T+T^{\prime}$.
@@
By Allen-Zhu & Li (2023) we already know Induction Hypothesis 15 holds at iteration $t=T$. Suppose Induction Hypothesis C.3 holds for all iterations $<t$ starting from $T$. We have established several lemmas:
@@
  • To prove 15(a), it suffices to plug (21), (22) into (18), use $\alpha_{p,v^{\prime}}\in[0,\gamma]$, use $|\mathcal{V}|=2k$, and use $|\langle w_{i,r}^{(t)},\xi_{p}\rangle|\leq\tilde{o}(\sigma_{0})$ from Lemma 31.
  • To prove 15(b), it suffices to plug (20), (21) into (18), use $\alpha_{p,v^{\prime}}\in[0,\gamma]$, use $|\mathcal{V}|=2k$, and use $|\langle w_{i,r}^{(t)},\xi_{p}\rangle|\leq\tilde{O}(\sigma_{0})$ from Lemma 31.
  • To prove 15(c), it suffices to plug (20), (21) into (19), use $\alpha_{p,v^{\prime}}\in[0,\gamma]$, use $|\mathcal{V}|=2k$, and use $|\langle w_{i,r}^{(t)},\xi_{p}\rangle|\leq\tilde{O}(\sigma_{0}\gamma k)$ from Lemma 31.
  • To prove 15(d), it suffices to note that (20) implies $\Phi_{i,l}^{(t)}\leq\tilde{O}(1)$, and note that Claim 24 implies $\Phi_{i,l}^{(t)}\geq\Omega(\Phi_{i,l}^{(0)})\geq\tilde{\Omega}(\sigma_{0})$.
  • To prove15(e), it suffices to invoke (22).
@@ E.3 Proof of Theorem 13
During the proof of Induction Hypothesis 15, i.e., the proofs of Lemma 29 to Lemma 32, we need the size of $\tilde{N}_{u}=|\tilde{Z}_{u}|$ larger than $\eta T^{\prime}\cdot\mathrm{poly}(k)$. As the probability that a sample $X\sim\mathcal{Z}_{u}$ after SA-CutOut belongs to $\tilde{\mathcal{Z}}_{u}$ is $1-\frac{1}{\mathrm{poly}(k)}$ (based on our definition of SA-CutOut and threshold $\tau$), we set the size of unlabeled dataset $N_{u}\geq\eta T^{\prime}\cdot\mathrm{poly}(k)$ in Parameter Assumption 10. Then recall our training objective is
@@
From Allen-Zhu & Li (2023) we know $\mathbb{E}_{(X,y)\sim\mathcal{Z}_{l}}[-\log\mathbf{logit}_{y}(F^{(t)},X)]\leq\frac{1}{\mathrm{poly}(k)}$ at the end of learning Phase I, and according to Claim 27 we now this continues to hold true during learning Phase II. For $\mathbb{E}_{X\sim\tilde{\mathcal{Z}}_{u}}[-\log\mathbf{logit}_{b}(F^{(t)},\mathcal{A}(X))]$, since we have for every data $(X,y)\sim\tilde{\mathcal{Z}}_{u}$ $(y=b)$:
@@
Therefore, by Claim 26, we know when $T^{\prime}\geq\frac{\mathrm{poly}(k)}{\eta}$,
$\frac{1}{T^{\prime}}\sum\nolimits_{t=T+T_{0}}^{T+T^{\prime}}\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u}}\left[-\log\mathbf{logit}_{y}(F^{(t)},X)\right]\leq\frac{1}{\mathrm{poly}(k)}.$
@@
$\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u}}[1-\mathbf{logit}_{y}(F^{(T+T^{\prime})},\mathcal{A}(X))]\leq\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u}}[-\log\mathbf{logit}_{y}(F^{(T+T^{\prime})},\mathcal{A}(X))]\leq\frac{1}{\mathrm{poly}(k)}.$
for the last iteration $T+T^{\prime}$, which directly implies that the training accuracy is perfect.
@@
$\Phi_{i}^{(T+T^{\prime})}-0.4\Phi_{j}^{(T+T^{\prime})}\geq\Omega(\log(k)),\quad\Phi_{i,1}^{(T+T^{\prime})},\Phi_{i,2}^{(T+T^{\prime})},\Phi_{j,1}^{(T+T^{\prime})},\Phi_{j,2}^{(T+T^{\prime})}\geq\Omega(\log(k)).$
@@ Appendix F Proof for FixMatch
In this section, we consider proving Theorem 11 on FixMatch. In this case, the formulation of strong augmentation $\mathcal{A}(\cdot)$ is defined in (10). For $(X,y)\in\tilde{Z}_{u,s}$ with $\hat{l}(X)=l^{*}$, the feature $v_{y,l^{*}}$ is masked with the probability $\pi_{2}$. When the patches of learned feature are masked, the left part is pure noise. In this way, same as Claim 23, for every $i\in[k]$, $\mathbf{logit}_{i}(F^{(t)},\mathcal{A}(X))=O(\frac{1}{k})$, and $V_{i,r,l}=\mathbb{I}_{v_{i,l}\in\mathcal{V}(X)}\sum_{p\in\mathcal{P}_{v_{i,l}}(X)}\overline{\operatorname{ReLU}}^{\prime}(\langle w_{i,r},x_{p}\rangle)z_{p}=o(\frac{1}{\mathrm{polylog}(k)})$. When the patches of noises are masked, the left part is semantic patches of feature $v_{y,l^{*}}$. Since we have already captured this feature in learning Phase I, we have $\mathbf{logit}_{y}(F^{(t)},\mathcal{A}(X))\geq 1-\tilde{O}(\frac{1}{s^{2}})$. In both above cases, by Claim 19 and Claim 20, the training samples $(X,y)\in\tilde{\mathcal{Z}}_{u,s}$ contributes little to the weight update process.
@@
For multi-view data $(X,y)\in\tilde{Z}_{u,m}$, when the patches of noises are masked with probability $1-\pi_{2}$, since the learned feature $v_{y,l^{*}}$ of Phase I is still in the data, we have $\mathbf{logit}_{y}(F^{(t)},\mathcal{A}(X))\geq 1-\tilde{O}(\frac{1}{s^{2}})$. When the patches of unlearned feature $v_{y,3-l^{*}}$ are masked with probability $\pi_{1}\pi_{2}$ or $(1-\pi_{1})\pi_{2}$ (depending on the value of $3-l^{*}$), the learned feature of Phase I is also still in the data. Thus we also have $\mathbf{logit}_{y}(F^{(t)},\mathcal{A}(X))\geq 1-\tilde{O}(\frac{1}{s^{2}})$. In this way, the loss on all data points $(X,y)\in\tilde{\mathcal{Z}}_{u,m}$ that belongs to the above two cases keeps small ($\leq\frac{1}{\mathrm{poly}(k)}$) and contributes negligible to the learning of unlearned features in Phase II. Finally, when the patches of learned feature $v_{y,l^{*}}$ are masked with probability $\pi_{1}\pi_{2}$ or $(1-\pi_{1})\pi_{2}$ (depending on the value of $l^{*}$), the remaining patches of feature $v_{y,3-l^{*}}$ are unlearned and the approximation of initial loss on this part of samples is the same as Claim 23. The loss on samples $(X,y)\sim\tilde{\mathcal{Z}}_{u,m}$ with learned features masked dominates the training objective in learning Phase II, and the rest proof schedule is the same as the proof of SA-FixMatch. However, now the size of data with learned features masked for class $i\in[k]$ is either $\pi_{1}\pi_{2}\cdot\tilde{N}_{u}^{i}$ or $(1-\pi_{1})\pi_{2}\cdot\tilde{N}_{u}^{i}$. Thus, similar to the proof of Theorem 13 and for the simplicity of notation, the requirement on the size of unlabeled data for FixMatch should be $\tilde{N}_{c}\geq\eta T^{\prime}\cdot\mathrm{poly}(k)/\min\{\pi_{1}\pi_{2},(1-\pi_{1})\pi_{2}\}$. Accordingly, we can derive that the relationship between the size of the unlabeled data in SA-FixMatch $N_{u}$ and FixMatch $N_{c}$ is given by $N_{u}=\max\{\pi_{1}\pi_{2},(1-\pi_{1})\pi_{2}\}N_{c}$.

### v2

@@ E.2.3 Noise Correlation is Small
Suppose Parameter 10 holds and suppose Induction Hypothesis 15 holds for all iteration $<t$ starting from $T_{1}$. For every $l\in[2]$, for every $r\in[m]$, for every $(X,y)\in\tilde{\mathcal{Z}}_{u}$ and $i\in[k]$:
@@
$\langle w_{i,r}^{(t+1)},\xi_{p}\rangle=\langle w_{i,r}^{(t)},\xi_{p}\rangle+\tilde{\Theta}(\frac{\eta}{\tilde{N}_{c}})\overline{\operatorname{ReLU}}^{\prime}(\langle w_{i,r}^{(t)},x_{p}\rangle)(1-\mathbf{logit}_{i}(F^{(t)},X))\pm\frac{\eta}{\sqrt{d}}.$
@@
$\langle w_{i,r}^{(t+1)},\xi_{p}\rangle=\langle w_{i,r}^{(t)},\xi_{p}\rangle-\tilde{\Theta}(\frac{\eta}{\tilde{N}_{c}})\overline{\operatorname{ReLU}}^{\prime}(\langle w_{i,r}^{(t)},x_{p}\rangle)\mathbf{logit}_{i}(F^{(t)},X)\pm\frac{\eta}{\sqrt{d}}.$
Suppose that it satisfies that $|\langle w_{i,r}^{(t)},x_{p}\rangle|\leq A$ for every $t<t_{0}$ where $t_{0}$ is any iteration $T_{1}\leq t_{0}\leq T_{1}+T_{2}$. When $T_{1}\leq t\leq T_{1}+T_{0}$, we have that
$\langle w_{i,r}^{(t)},\xi_{p}\rangle$ $\leq\langle w_{i,r}^{(T_{1})},\xi_{p}\rangle+\tilde{O}\left(\frac{T_{0}\eta A^{q-1}}{\tilde{N}_{c}}\right)+\frac{T_{0}\eta}{\sqrt{d}}$
$\leq\tilde{o}(\sigma_{0})+\tilde{O}\left(\frac{kA^{q-1}}{\tilde{N}_{c}\sigma_{0}^{q-2}}\right)+\frac{T_{0}\eta}{\sqrt{d}},$
where the last step is because $T_{0}=\tilde{\Theta}(\frac{k}{\eta\sigma_{0}^{q-2}})$. When $T_{1}+T_{0}\leq t\leq T_{1}+T_{2}$, for multi-view data $(X,y)\sim\tilde{\mathcal{Z}}_{u,m}$, based on (15) in Claim 28, we can obtain that
$\langle w_{i,r}^{(t)},\xi_{p}\rangle$ $\leq\langle w_{i,r}^{(T_{1}+T_{0})},\xi_{p}\rangle+\tilde{O}\left(\frac{k^{5}A^{q-1}}{s^{2}\tilde{N}_{c}}\right)+\frac{(T_{2}-T_{0})\eta}{\sqrt{d}}$
$\leq\tilde{O}\left(\frac{kA^{q-1}}{\tilde{N}_{c}\sigma_{0}^{q-2}}+\frac{k^{5}A^{q-1}}{s^{2}\tilde{N}_{c}}\right)+\frac{T_{2}\eta}{\sqrt{d}}.$
@@
$\langle w_{i,r}^{(t)},\xi_{p}\rangle$ $\leq\langle w_{i,r}^{(T_{1}+T_{0})},\xi_{p}\rangle+\tilde{O}\left(\frac{T_{2}\eta A^{q-1}}{\tilde{N}_{c}}\right)+\frac{(T_{2}-T_{0})\eta}{\sqrt{d}}$
$\leq\tilde{O}\left(\frac{kA^{q-1}}{\tilde{N}_{c}\sigma_{0}^{q-2}}+\frac{T_{2}\eta A^{q-1}}{\tilde{N}_{c}}\right)+\frac{T_{2}\eta}{\sqrt{d}}.$
When $p\in\mathcal{P}_{v_{i,l}}(X)$, we have $|\langle w_{i,r}^{(t)},x_{p}\rangle|\leq\tilde{O}(1)$ from Induction Hypothesis 15. Then plugging in $A=\tilde{O}(1),\tilde{N}_{c}\geq\tilde{\Omega}\left(\frac{k}{\sigma_{0}^{q-1}}\right)$, $\tilde{N}_{c}\geq\tilde{\Omega}\left(\frac{k^{5}}{\sigma_{0}}\right)$ and $\tilde{N}_{c}\geq\eta T_{2}\mathrm{poly}(k)$, we can obtain that $|\langle w_{i,r}^{(t)},\xi_{p}\rangle|\leq\tilde{o}(\sigma_{0})$.
When $p\in\mathcal{P}(X)\setminus(\mathcal{P}_{v_{i,1}}(X)\cup\mathcal{P}_{v_{i,2}}(X))$, we have $|\langle w_{i,r}^{(t)},x_{p}\rangle|\leq\tilde{O}(\sigma_{0})$ from the Induction Hypothesis 15. Then plugging in $A=\tilde{O}(\sigma_{0}),\tilde{N}_{c}\geq k^{5}$ and $\tilde{N}_{c}\geq\eta T_{2}\mathrm{poly}(k)$, we can obtain that $|\langle w_{i,r}^{(t)},\xi_{p}\rangle|\leq\tilde{O}(\sigma_{0})$.
When $p\in[P]\setminus\mathcal{P}(X)$, we have $|\langle w_{i,r}^{(t)},x_{p}\rangle|\leq\tilde{O}(\sigma_{0}\gamma k)$ from Induction Hypothesis 15. Then plugging in $A=\tilde{O}(\sigma_{0}\gamma k),\tilde{N}_{c}\geq k^{5}$ and $\tilde{N}_{c}\geq\eta T_{2}\mathrm{poly}(k)$, we can obtain that $|\langle w_{i,r}^{(t)},\xi_{p}\rangle|\leq\tilde{O}(\sigma_{0}\gamma k)$.
@@ E.2.4 Diagonal Correlations are Nearly Non-Negative
Suppose 10 holds and suppose Induction Hypothesis 15 holds for all iteration $<t$ starting from $T_{1}$. Then,
@@
From Lemma D.27 in Allen-Zhu & Li (2023), we know $\langle w_{i,r}^{(t)},v_{i,l}\rangle\geq-\tilde{O}(\sigma_{0})$ for every iteration $t\leq T_{1}$. Now we consider any iteration $t>T_{1}$ so that $\langle w_{i,r}^{(t)},v_{i,l}\rangle\leq-\widetilde{\Omega}(\sigma_{0})$. We start from this iteration to see how negative the next iterations can be. Without loss of generality, we consider the case when $\langle w_{i,r}^{(t^{\prime})},v_{i,l}\rangle\leq-\widetilde{\Omega}(\sigma_{0})$ holds for every $t^{\prime}\geq t$. By Claim 19 and Claim 20,
@@
We first consider every $t\leq T_{1}+T_{0}$. Using Claim 23 we have $\mathbf{logit}_{i}(F^{(t)},X)=O\left(\frac{1}{k}\right)$, which implies
@@
As for $t>T_{1}+T_{0}$, combining with Claim 26 and the fact that $\mathbf{logit}_{i}(F^{(t)},X)\leq 1-\mathbf{logit}_{y}(F^{(t)},X)$ for $i\neq y$, we have
@@ E.2.5 Proof of Induction Hypothesis 15
Under Parameter Assumption 10, for any $m=\mathrm{polylog}(k)$ and sufficiently small $\eta\leq\frac{1}{\mathrm{poly}(k)}$, our Induction Hypothesis 15 holds for all iterations $t=T_{1},T_{1}+1,\dots,T_{1}+T_{2}$.
@@
By Allen-Zhu & Li (2023) we already know Induction Hypothesis 15 holds at iteration $t=T_{1}$. Suppose Induction Hypothesis 15 holds for all iterations $<t$ starting from $T_{1}$. We have established several lemmas:
@@
  • To prove 15(a), it suffices to plug Eq. (21), Eq. (22) into Eq. (18), use $\alpha_{p,v^{\prime}}\in[0,\gamma]$, use $|\mathcal{V}|=2k$, and use $|\langle w_{i,r}^{(t)},\xi_{p}\rangle|\leq\tilde{o}(\sigma_{0})$ from Lemma 31.
  • To prove 15(b), it suffices to plug Eq. (20), Eq. (21) into Eq. (18), use $\alpha_{p,v^{\prime}}\in[0,\gamma]$, use $|\mathcal{V}|=2k$, and use $|\langle w_{i,r}^{(t)},\xi_{p}\rangle|\leq\tilde{O}(\sigma_{0})$ from Lemma 31.
  • To prove 15(c), it suffices to plug Eq. (20), Eq. (21) into Eq. (19), use $\alpha_{p,v^{\prime}}\in[0,\gamma]$, use $|\mathcal{V}|=2k$, and use $|\langle w_{i,r}^{(t)},\xi_{p}\rangle|\leq\tilde{O}(\sigma_{0}\gamma k)$ from Lemma 31.
  • To prove 15(d), it suffices to note that Eq. (20) implies $\Phi_{i,l}^{(t)}\leq\tilde{O}(1)$, and note that Claim 24 implies $\Phi_{i,l}^{(t)}\geq\Omega(\Phi_{i,l}^{(0)})\geq\tilde{\Omega}(\sigma_{0})$.
  • To prove 15(e), it suffices to invoke Eq. (22).
@@ E.3 Proof of Theorem 13
Recall our training objective is
@@
From Allen-Zhu & Li (2023) we know $\mathbb{E}_{(X,y)\sim\mathcal{Z}_{l}}[-\log\mathbf{logit}_{y}(F^{(T_{1})},X)]\leq\frac{1}{\mathrm{poly}(k)}$ holds at the end of learning Phase I, and according to Claim 27 we now this continues to hold true during learning Phase II. For $\mathbb{E}_{X\sim\tilde{\mathcal{Z}}_{u}}[-\log\mathbf{logit}_{b}(F^{(t)},\mathcal{A}(X))]$, since we have for every data $(X,y)\sim\tilde{\mathcal{Z}}_{u}$ $(y=b)$:
@@
Therefore, by Claim 26, we know when $T_{2}\geq\frac{\mathrm{poly}(k)}{\eta}$,
$\frac{1}{T_{2}}\sum\nolimits_{t=T_{1}+T_{0}}^{T_{1}+T_{2}}\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u}}\left[-\log\mathbf{logit}_{y}(F^{(t)},X)\right]\leq\frac{1}{\mathrm{poly}(k)}.$
@@
$\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u}}[1-\mathbf{logit}_{y}(F^{(T)},\mathcal{A}(X))]\leq\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u}}[-\log\mathbf{logit}_{y}(F^{(T)},\mathcal{A}(X))]\leq\frac{1}{\mathrm{poly}(k)}$
for the last iteration $T=T_{1}+T_{2}$, which directly implies that the training accuracy is perfect.
@@
$\Phi_{i}^{(T)}-0.4\Phi_{j}^{(T)}\geq\Omega(\log(k)),\quad\Phi_{i,1}^{(T)},\Phi_{i,2}^{(T)},\Phi_{j,1}^{(T)},\Phi_{j,2}^{(T)}\geq\Omega(\log(k)).$
@@ Appendix F Proof for FixMatch
In this section, we consider proving Theorem 11 on FixMatch. In this case, the formulation of strong augmentation $\mathcal{A}(\cdot)$ is defined in Eq. (10). For $(X,y)\in\tilde{Z}_{u,s}$ with $\hat{l}(X)=l^{*}$, the feature $v_{y,l^{*}}$ is removed with the probability $\pi_{2}$. When the patches of learned feature are removed, the left part is pure noise. In this way, same as Claim 23, for every $i\in[k]$, $\mathbf{logit}_{i}(F^{(t)},\mathcal{A}(X))=O(\frac{1}{k})$, and $V_{i,r,l}=\mathbb{I}_{v_{i,l}\in\mathcal{V}(X)}\sum_{p\in\mathcal{P}_{v_{i,l}}(X)}\overline{\operatorname{ReLU}}^{\prime}(\langle w_{i,r},x_{p}\rangle)z_{p}=o(\frac{1}{\mathrm{polylog}(k)})$. When the patches of noises are removed, the left part is semantic patches of feature $v_{y,l^{*}}$. Since we have already captured this feature in learning Phase I, we have $\mathbf{logit}_{y}(F^{(t)},\mathcal{A}(X))\geq 1-\tilde{O}(\frac{1}{s^{2}})$. In both above cases, by Claim 19 and Claim 20, the training samples $(X,y)\in\tilde{\mathcal{Z}}_{u,s}$ contribute little to the weight update process.
@@
For multi-view data $(X,y)\in\tilde{Z}_{u,m}$, when the patches of noises are removed with probability $1-\pi_{2}$, since the learned feature $v_{y,l^{*}}$ of Phase I is still in the data, we have $\mathbf{logit}_{y}(F^{(t)},\mathcal{A}(X))\geq 1-\tilde{O}(\frac{1}{s^{2}})$. When the patches of unlearned feature $v_{y,3-l^{*}}$ are removed with probability $\pi_{1}\pi_{2}$ or $(1-\pi_{1})\pi_{2}$ (depending on the value of $3-l^{*}$), the learned feature of Phase I is also still in the data. Thus we also have $\mathbf{logit}_{y}(F^{(t)},\mathcal{A}(X))\geq 1-\tilde{O}(\frac{1}{s^{2}})$. In this way, the loss on all data points $(X,y)\in\tilde{\mathcal{Z}}_{u,m}$ that belongs to the above two cases keeps small ($\leq\frac{1}{\mathrm{poly}(k)}$) and contributes negligible to the learning of unlearned features in Phase II. Finally, when the patches of learned feature $v_{y,l^{*}}$ are removed with probability $\pi_{1}\pi_{2}$ or $(1-\pi_{1})\pi_{2}$ (depending on the value of $l^{*}$), the remaining patches of feature $v_{y,3-l^{*}}$ are unlearned and the approximation of initial loss on this part of samples is the same as Claim 23. The loss on samples $(X,y)\sim\tilde{\mathcal{Z}}_{u,m}$ with learned features removed dominates the training objective in learning Phase II, and the rest proof schedule is the same as the proof of SA-FixMatch. However, now the size of data with learned features removed for class $i\in[k]$ is either $\pi_{1}\pi_{2}\cdot\tilde{N}_{u}^{i}$ or $(1-\pi_{1})\pi_{2}\cdot\tilde{N}_{u}^{i}$. Thus, similar to the proof of Theorem 13 and for the simplicity of notation, the requirement on the size of unlabeled data for FixMatch should be $\tilde{N}_{u}\geq\eta T_{2}\cdot\mathrm{poly}(k)/\min\{\pi_{1}\pi_{2},(1-\pi_{1})\pi_{2}\}$. Accordingly, we can derive that the relationship between the size of the unlabeled data in SA-FixMatch $N_{c}$ and FixMatch $N_{u}$ is given by $N_{c}=\max\{\pi_{1}\pi_{2},(1-\pi_{1})\pi_{2}\}N_{u}$.